OpenStudy (anonymous):

Solve the following formula for the specified variable. C=1/4h(s+k) solve for s???

OpenStudy (anonymous):

if this was a normal equation, and all those variables were numbers, how would you solve for s? hint: try to get s by itself

OpenStudy (anonymous):

It's been so long since i actually did this type of problem

OpenStudy (anonymous):

i also have a question about the original question is it C=(1/4)h(s+k) or C=1/(4h(s+k))

OpenStudy (anonymous):

it's the first one that you put on there

OpenStudy (anonymous):

c=1/4h(s+k)

OpenStudy (anonymous):

the first and easiest thing to get rid of is the 1/4 if we want to re-write the equation, we could like this: c=(h(s+k))/4 since we are dividing by 4 on the right, we can multiply both sides by 4 to get closer to having s by itself (by moving the 4 over) we get: 4c=h(s+k) now, to get rid of that h notice its being multiplied by the quantity inside the parenthesis to move it to the other side, we need to divide both sides by h (4c)/h=s+k finally, subtract k from both sides to get your answer

OpenStudy (anonymous):

ok im confused

OpenStudy (anonymous):